# Transferring Textual Preferences to Vision-Language Understanding through Model Merging

### v1

@@ Abstract
Large vision-language models (LVLMs) perform outstandingly across various multimodal tasks. However, their ability to evaluate generated content remains limited, and training vision-language reward models (VLRMs) with preference data is computationally expensive. This paper explores a training-free alternative by merging text-based reward models (RMs) with LVLMs to create VLRMs. Our approach shows that integrating these models leads to improved performance over LVLMs’ scoring and text-based RMs, offering an efficient method for incorporating textual preferences into LVLMs.
@@ 1 Introduction
Figure 1: Framework for merging a text-based RM with an LVLM. LVLMs excel at visual tasks, while text-based RMs struggle to provide accurate rewards without visual cues. We transfer textual preferences to the vision-language understanding, resulting in a VLRM.
@@ 3.1 Model Components
$\theta^{\text{PRE}}=\{\theta^{\text{PRE}}_{\text{emb}},\theta^{\text{PRE}}_{\text{trans}},\theta^{\text{PRE}}_{\text{lm}}\},\vspace{-3pt}$
@@
$\theta^{\text{LVLM}}=\{\theta^{\text{LVLM}}_{\text{venc}},\theta^{\text{LVLM}}_{\text{adapt}},\theta^{\text{LVLM}}_{\text{emb}},\theta^{\text{LVLM}}_{\text{trans}},\theta^{\text{LVLM}}_{\text{lm}}\},\vspace{-3pt}$
@@
$\theta^{\text{RM}}=\{\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{rm}}\},\vspace{-3pt}$
@@ 3.2 Merging Strategies
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot\theta^{\text{LVLM}}_{\text{trans}}+(1-\lambda)\cdot\theta^{\text{RM}}_{\text{trans}},\vspace{-3pt}$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot\tau_{\text{LVLM}}+\lambda\cdot\tau_{\text{RM}},$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),\vspace{-3pt}$
@@ 3.3 Merged VLRM
$\theta^{\text{MERGE}}=\{\theta^{\text{LVLM}}_{\text{venc}},\theta^{\text{LVLM}}_{\text{adapt}},\theta^{\text{MERGE}}_{\text{emb}},\theta^{\text{MERGE}}_{\text{trans}},\theta^{\text{RM}}_{\text{rm}}\},\vspace{-3pt}$
@@ 4.3 Main Results
As shown in Table 1, merged VLRMs consistently outperform Llama-3.2-Vision and Tulu-2.5-RM across nearly all merging methods and benchmarks. This result demonstrates that combining a text-based RM with an LVLM effectively transfers textual preferences without training. Different merging strategies achieve the highest scores in different benchmarks, but overall, more advanced methods outperform simpler ones, highlighting the advantages of structured merging techniques. Additionally, in several benchmarks, merged VLRMs surpass or match the strong Cascade baseline, suggesting that model merging captures more information than merely cascading two models. Notably, in VL-RewardBench, our merged VLRMs even exceed the performance of the 90B LVLM and achieve results comparable to commercial models. A similar trend emerges when using Tulu-3-RM as the text-based RM; further details are provided in Appendix G.1.
@@ Appendix A Merging Details
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot\tau_{\text{LVLM}}+\lambda\cdot\tau_{\text{RM}},$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),$
@@ G.3 Effect of Merging Hyperparameters
In this study, we optimize hyperparameter merging using sampled instances from RLAIF-V. The results, based on 400 sampled RLAIF-V instances used as a validation set, are presented in Tables 21 to 21. Bold text highlights the best performance, while text with * indicates cases where scores are tied. In these cases, an additional 100 samples are used, and * marks the top-performing result among them.

### v2

@@ Abstract
Large vision-language models (LVLMs) perform outstandingly across various multimodal tasks. However, their ability to evaluate generated content remains limited, and training vision-language reward models (VLRMs) with preference data is computationally expensive. This paper explores a training-free alternative by merging text-based reward models (RMs) with LVLMs to create VLRMs. Our approach shows that integrating these models leads to improved performance over LVLMs’ scoring and text-based RMs, offering an efficient method for incorporating textual preferences into LVLMs. The code and data are publicly available at https://github.com/lca0503/MergeToVLRM.
@@ 1 Introduction
Figure 1: Framework for merging a text-based RM with an LVLM. LVLMs excel at visual tasks, while text-based RMs struggle to provide accurate rewards without visual cues. We transfer textual preferences to the vision-language understanding, resulting in a VLRM. All icons used in this figure are sourced from https://www.flaticon.com/
@@ 3.1 Model Components
$\theta^{\text{PRE}}=\{\theta^{\text{PRE}}_{\text{emb}},\theta^{\text{PRE}}_{\text{trans}},\theta^{\text{PRE}}_{\text{lm}}\},$
@@
$\theta^{\text{LVLM}}=\{\theta^{\text{LVLM}}_{\text{venc}},\theta^{\text{LVLM}}_{\text{adapt}},\theta^{\text{LVLM}}_{\text{emb}},\theta^{\text{LVLM}}_{\text{trans}},\theta^{\text{LVLM}}_{\text{lm}}\},$
@@
$\theta^{\text{RM}}=\{\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{rm}}\},$
@@ 3.2 Merging Strategies
$\theta^{\text{MERGE}}_{\text{trans}}=\theta^{\text{PRE}}_{\text{trans}}+\lambda\cdot\tau_{\text{LVLM}}+\lambda\cdot\tau_{\text{RM}},$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\theta^{\text{PRE}}_{\text{trans}}+\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),$
@@ 3.3 Merged VLRM
$\theta^{\text{MERGE}}=\{\theta^{\text{LVLM}}_{\text{venc}},\theta^{\text{LVLM}}_{\text{adapt}},\theta^{\text{MERGE}}_{\text{emb}},\theta^{\text{MERGE}}_{\text{trans}},\theta^{\text{RM}}_{\text{rm}}\},$
@@ 4.3 Main Results
As shown in Table 1, merged VLRMs consistently outperform Llama-3.2-Vision and Tulu-2.5-RM across nearly all merging methods and benchmarks. This result demonstrates that combining a text-based RM with an LVLM effectively transfers textual preferences without training. Different merging strategies achieve the highest scores in different benchmarks, but overall, more advanced methods outperform simpler ones, highlighting the advantages of structured merging techniques. Additionally, in several benchmarks, merged VLRMs surpass or match the strong Cascade baseline, suggesting that model merging captures more information than merely cascading two models. Furthermore, as shown in Table 2, our merged VLRMs even exceed the performance of the 90B LVLM and achieve results comparable to commercial models. A similar trend emerges when using Tulu-3-RM as the text-based RM; further details are provided in Appendix G.1.
Method General Hallucination Reasoning
Open-Source Models*
Llama-3.2-Vision (11B) 33.3 38.4 56.6
Llama-3.2-Vision (90B) 42.6 57.3 61.7
Proprietary Models*
Gemini-1.5-Flash 47.8 59.6 58.4
Gemini-1.5-Pro 50.8 72.5 64.2
GPT-4o-mini 41.7 34.5 58.2
GPT-4o 49.1 67.6 70.5
Using TULU-2.5-RM for merging
Linear 39.3 52.3 54.4
Task Vec. 48.6 59.4 59.7
TIES 43.7 58.2 58.5
DARE + Task Vec. 49.2 61.7 61.0
DARE + TIES 49.2 59.1 58.2
Table 2: VL-RewardBench results comparing open-source and proprietary models with our reward model using TULU-2.5-RM for merging. *Indicates results from Li et al. (2024a). Full results are provided in Table 12
@@ 4.4 Analysis
Computation Overhead
In our experiments, model merging is done entirely on CPUs (Intel Xeon Silver 4216) using a system with 128 GB of RAM. Using 11 different $\lambda$ values for weighted averaging takes about 1.5 hours of CPU time. The task arithmetic method takes a similar amount of time when using the same number of $\lambda$ values. Applying 12 combinations of $\lambda$ and density $d$ for the TIES method takes about 6 hours of CPU time, while DARE takes around 3 hours to handle the same number of combinations.
We evaluate the models on a validation set of 400 examples from the RLAIF-V dataset. We run model inference on GPUs with 24 GB of memory (Nvidia GeForce RTX 3090). Across all configurations and merging methods, inference takes approximately 1.5 hours of GPU time per method.
Overall, merging and evaluation require much less computing time than training a reward model from scratch. Since merging is the most time-consuming step and runs only on the CPU, the total computational cost stays relatively low. Also, both merging and evaluation can be run in parallel on multiple machines to reduce the actual runtime.
@@ 5 Conclusion
Acknowledgements
We thank the reviewers for their insightful comments. This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grant 113-2628-E-002-033. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources. We are also grateful to Yu-Xiang Lin from National Taiwan University for his valuable advice and thoughtful discussions.
@@ Appendix A Merging Details
$\theta^{\text{MERGE}}_{\text{trans}}=\theta^{\text{PRE}}_{\text{trans}}+\lambda\cdot\tau_{\text{LVLM}}+\lambda\cdot\tau_{\text{RM}},$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\theta^{\text{PRE}}_{\text{trans}}+\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),$
@@
$\theta^{\text{MERGE}}_{\text{trans}}=\theta^{\text{PRE}}_{\text{trans}}+\lambda\cdot f(\tau^{\text{LVLM}},d)+\lambda\cdot f(\tau^{\text{RM}},d),$
@@ G.3 Effect of Merging Hyperparameters
In this study, we optimize hyperparameter merging using sampled instances from RLAIF-V. The results, based on 400 sampled RLAIF-V instances used as a validation set, are presented in Tables 22 to 22. Bold text highlights the best performance, while text with * indicates cases where scores are tied. In these cases, an additional 100 samples are used, and * marks the top-performing result among them.